OpenStudy (anonymous):

If f(x) = x + 4, find f(3).

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Thanks!(:

OpenStudy (shamim):

given f(x)=x+4

OpenStudy (shamim):

so f(3)=3+4=7